OpenStudy (rva.raghav):

The circuit parameters have the following values: VS=4.5V and R=1kΩ, and the ExpFET parameters are: K=1mA, VT=1V and VN=1V. Given that vIN=0.5V, what is the value of vOUT in Volts? unanswered Given that vIN=2V, what is the value of vOUT in Volts? unanswered What is the value of vIN in Volts when vOUT drops to VS2? unanswered Now you replace the ExpFET with a MOSFET with parameters K=1mAV2, and VT=1V. Assuming the MOSFET is operating in the saturation region, what is the value of vIN in volts when vOUT drops to VS2? (Recall that for a MOSFET in saturation: iD=K2(VGS−VT)2

OpenStudy (anonymous):

You'll find the formulas here: http://openstudy.com/study#/updates/50d36174e4b0b19ec2190832

OpenStudy (rva.raghav):

i want answers

OpenStudy (rva.raghav):

please anyone

OpenStudy (anonymous):

Nobody is going to give you answers. If you can't even put values in a formula, then there's no point in passing this exam.

OpenStudy (rva.raghav):

just shetup and getlost

OpenStudy (anonymous):

Haha, lost your brain?

OpenStudy (rva.raghav):

the pronlem1 formula of q2 i cant understand

OpenStudy (rva.raghav):

sry rmpk

OpenStudy (anonymous):

K,UTCFGRJYMB,Y

OpenStudy (nurali):

@rva.raghav 1. 4.5 2. 2.7818 3. 2.1786 4. 3.1213 5. -2.7182 6. -1

OpenStudy (nurali):

@rva.raghav if correct please click best response.

OpenStudy (anonymous):

Q1 a)ia=2*i1*R/(2*R+rm) and vo=ia*rm b)iA =(-i2)*R/(2*R+rm) and V0=ia*rm c) part a) + part b) d)2R Q2 p1: Vs as vin<vt p2: v0 = vs-kr(e^((vin-vt)/vn)-1) p3: vin = ln(vs/2kr +1)vn+vt p4: vin=sqrt(vs/kr)+vt p5: -kre^((vin-vt)/vn) p6: -kr(vin-vt) Q3 same as HP8p1 Q4 p1=V*R2/(R1+R2)*(1-e^(-(R1+R2)*t/(R1*R2*C))) p2=(V*R2/(R1+R2))*cos(1/sqrt(L*C)*(t-T)) p3=C*(V*R2/(R1+R2))*sin((t-T)*1/sqrt(L*C))/sqrt(L*C) Q5 E V B Y F Z

OpenStudy (anonymous):

@rishikesh.147 If you're so smart, give the formula for Q6. Wait, you can't, cause all you do is copy other's formulas, trying to get credit for them. Haha, you're so lame!